Shift report 20010914 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Sep 14 2001 - 00:04:11 EDT

  • Next message: Apache: "Shift report 20010914 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Sep 14 00:04:11 2001, Shift: 16:00-24:00
    Supervisor: R. Debbe, with: none whatsoever !!
    Log book#8 pages 123 - 127
    
    Shift Summary:
    ------------------------------------------------------------
    Machine stat: At begining of shift the machine was on RF studies, we took data 
    and can see the vertex with a narrow component (a nipple over the standard peak) at 18:30
    they ended the studies and went back to 56 bunches and lower frequency on RF.
    in 10 minutes the beam was ready for physics!! the same store 
    is still going with some 60 ZDC coinc. (it started with 300).
    2 Achievements: I only ate two donuts and 10 chocolates. Meanwhile
    the daq collected the following runs:
    4717 200 MHz RF magnets off
    4720 data with 374K trigg 1
    4721 data with 134K trigg 1
    4723 still going
    
    3 Detector problems: I looked at T3 and found the field wires
    with current that fluctuates from 0 to 700 nA without beam,
    when the beam was present at the highest intensities the voltage
    fluctuates by as much as 30 volts.
    The current in T3A1 was above 100 microA, FV adviced to lower
    40 volts. After that the current was ~80 microA. T5A1 has also currents
    above 100 microA. At end of shift T5 anodes tripped and a few minutes
    later T4A1 tripped. No possible explanation.
    
    I retimed the RICH: delayed the gate by 60 nsec now the data
    should be cleaner.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Sep 14 2001 - 00:04:45 EDT